HUBBLE SPACE TELESCOPE DAILY REPORT #3100 PERIOD COVERED: 0000Z (UTC) 04/22/02 - 0000Z (UTC) 04/23/02 Daily Status Report as of 113/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed Three Sets of ACS/WFC/HRC 9562 (Cycle 11 Internal Flat Field Stability) The Advance Camera for Surveys (WFC and HRC) was used to obtain flat field stability and characterization samples of the ACS filter set. Only internal exposures using the calibration lamps were required. There were no reported problems. 1.2 Completed FGS/1 9034 (The Masses and Luminosities of Population II Stars) Fine Guidance Sensor #1R was used to observe the mass-luminosity relation {MLR} of Population II stars of which very little is currently known. With the advent of the Hipparcos Catalogue, improved distances to many spectroscopic binaries known to be Pop II systems are now available. After surveying the literature and making reasonable estimates of the secondary masses, we find 13 systems whose minimum separation should be larger than the resolution limit of FGS #1. The observations completed nominally. 1.3 Completed Four Sets of WF/PC-2 9318 (POMS Test Proposal: WFII Parallel Archive Proposal Continuation) The WF/PC-2 was used to perform the generic target version of the WFPC2 Archival Pure Parallel program. The program was used to take parallel images of random areas of the sky, following the recommendations of the Parallels Working Group. There were no reported problems. 1.4 Completed STIS/CCD 8907 (Spectroscopic Flats C10) The Space Telescope Imaging Spectrograph (CCD) was used to obtain CCD flats in the spectrographic mode. There were no reported problems. 1.5 Completed Five Sets of ACS/WFC/HRC 8947 (Weekly Test) The Advanced Camera for Surveys (WFC and HRC) was used to perform basic tests to monitor, the read noise, the development of hot pixels and test for any source of noise in ACS CCD detectors. This program will be executed at least once a day for the entire lifetime of ACS. There were no problems reported. 1.6 Completed Four Sets of WF/PC-2 8937 (Cycle 9 Supplemental Darks pt2/3) The WF/PC-2 was used obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. No problems were encountered. 1.7 Completed STIS/CCD/MA1 9094 (The Wind-Wind Interactions in HD 5980) The Space Telescope Imaging Spectrograph CCD and MA1) was used to observe HD 5980, the visually brightest object in the SMC. It contains a close eclipsing binary LBV+WR {=Star A+Star B}, the only one known to exist, with the possible exception of Eta Carinae. The LBV underwent a great eruption in 1994, ejecting <10^-3 M_odot at velocities in the range of 200-800 km s^-1. After the eruption, its wind velocity has been steadily increasing, and producing a circumstellar {CSM} fast- wind/slow-wind interaction region, which has now become visible spectroscopically in the C IV and Si IV lines at an expansion velocity of -680 km s^-1. These narrow features varied over a 1 year timescale. It is likely that HD 5980 is forming an H II region analogous to the Homunculus in Eta Carinae. In addition, the wind-wind collision shocked region between the two stars currently {1999-2000} forms a shock cone that winds tightly around the WR star. The purpose of this proposal is to: a} Determine the LBV's wind velocity close to the time when it will be observed with XMM and CHANDRA; b} follow the changes in the high velocity CSM absorption features; and c} Constrain the radial velocity variability of the third component, Star C, in order to determine whether it is bound to the A+B pair. All observations completed without incident. 1.8 Completed Thirteen Sets of NICMOS/1/2/3 8945 (Cooling System Monitoring) The Near Infrared Camera and Multi-Object Camera (NIC1, NIC2 and NIC3 was used 1) to measure NICMOS detector performance during the cool-down and steady state operation of the NCS. 2}. to demonstrate stability {+/-0.1K} of the NICMOS detector temperature at the optimal science operating temperature. and 3}.to demonstrate repeatability {+/-0.1 K} of NICMOS detector temperature following changes from the optimal science operating temperature. No problems were encountered. 1.9 Completed Five Sets of WF/PC-2 8935 (Cycle 10 Standard Darks) The WF/PC-2 was used to obtain dark frames every week in order to provide data for the ongoing calibration of the CCD dark current rate, and to monitor and characterize the evolution of hot pixels. Over an extended period these data will also provide a monitor of radiation damage to the CCDs. The proposal completed with no reported problems. 1.10 Completed Two Sets of STIS/MA1 8920 (Cycle 10 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA1) was used to perform the routine monitoring of the MAMA detector dark noise, and is the primary means of checking on health of the MAMA detectors systems through frequent monitoring of the background count rate. The proposal completed with no reported anomalous activity. 1.11 Completed Three Sets of STIS/CCD 8902 (Dark Monitor-Part 2) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the darks. No problems were reported. 1.12 Completed Two Sets of STIS/CCD/MA2 9496 (UV Spectroscopic Investigation of Any Bright, Newly Discovered Comet) The Space Telescope Imaging Spectrograph (CCD and MA1) was used to perform a Target of Opportunity program to investigate any bright comet {mboxV<=sssim7} that is newly discovered during Cycle 11, including comets of any dynamical class. The main scientific objective is to obtain accurate abundance measurements for several key cometary species: CO from the CO 4PG bands, cotwo from the CO Cameron bands, stwo from the stwo B-X bands, cstwo from CS emissions, and water from OH emissions. The UV Cameron band emission is currently the only way to probe cotwo in comets. No anomalous activity occurred. 1.13 Completed WF/PC-2 8939 (Cycle 10 Internal Monitor) The WF/PC-2 was used to calibrate the internal monitor, to be run weekly to monitor the health of the cameras. No problems were encountered. 1.14 Completed ACS/WFC 9575 (Default {Archival} Pure Parallel Program) The Advanced Camera for Surveys (WFC) was used to test ACS pure parallels in POMS. There were no reported problems. 1.15 Completed FGS/1 9581 (A BEA Science Program: A High Angular Resolution Survey of Some of the Galaxy's Most Massive Stars) Fine Guidance Sensor #1R was used to make optimal use of otherwise unused telescope time while HST's pointing is restricted to the BEA cone during the first 12 days of SMOV3B. The launch date of Feb 29 placed two of the Galaxy's most active star forming regions, Trumplers 14 and 16 in the Carina Nebula, within the BEA for the entire 12 day period. FGS1r was used to conduct a high angular resolution survey of the most massive and luminous stars in these two magnificent clusters to establish the incidence of binarity at the high mass end of the HR diagram. This survey will close the detection gap between spectroscopic and speckle surveys in which duplicity can be concealed and will dramatically improve our knowledge of the IMF at the high mass end of the main sequence as well as the luminosity function {LF} of star forming regions which produce such high mass stars. Finally, this survey will identify targets for an approved SIM Key Science project to dynamically determine O-star masses. There was no anomalous activity. 1.16 Completed STIS/MA2 9573 (NUV-MAMA Daily Dark Monitor) Abstract The Space Telescope Imaging Spectrograph (CCD and MA2) was used to perform daily monitoring of the NUV MAMA detector dark noise, monitoring the effects of thermal changes on the NUV dark rate. No anomalous activity was noted. 1.17 Completed STIS/CCD 8904 (Bias Monitor-Part 2) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the bias in the 1x1, 1x2, 2x1, and 2x2 bin settings at gain=1, and 1x1 at gain = 4 in order to build up high-S/N superbiases and track the evolution of hot columns. The proposal completed with no anomalous activity. 1.18 Completed Three Sets of STIS/CCD 9066 (Closing in on the Hydrogen Reionization Edge of the Universe) The Space Telescope Imaging Spectrograph (CCD) was used in parallel constrain the Hydrogen reionization edge in emission that marks the transition from a neutral to a fully ionized IGM at a predicted redshifts. The proposal completed uneventfully. 1.19 Completed Three Sets of ACS/WFC/HRC 9075 (Cosmological Parameters from Type Ia Supernovae at High Redshift) The Advanced Camera for Surveys (WFC and HRC) was used to obtain a Hubble diagram of Type Ia supernovae {SNe Ia} that will be of long lasting value as a record of the expansion history of the universe. No problems were encountered. 1.20 Completed STIS/CCD 9131 (Imaging the Host Galaxies of High Redshift Type Ia Supernovae) The Space Telescope Imaging Spectrograph (CCD) was used to complete the snapshot survey of distant galaxies of known redshift which hosted supernovae {SNe} of Type Ia found via the Supernova Cosmology Project {SCP}. No problems were seen. 1.21 Completed WF/PC-2 9319 (POMS Test Proposal: WFII Backup Parallel Archive Proposal II) The WF/PC-2 was used to execute a POMS test proposal, designed to simulate future scientific plans. The proposal completed without incident. 1.22 Completed STIS/CCD 8916 (Spectroscopic Sensitivity Workout: First-order Modes) The Space Telescope Imaging Spectrograph (CCD) was used to measure the basic sensitivity for all supported MAMA and CCD first-order spectroscopic modes. Sensitivity measurements are done for all supported tilts of the gratings, at a S/N suitable to any particular setting, in order to get all measurements done in a reasonable number of orbits but still get a very accurate sensitivity measurement. There were no reported problems. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 11 Successful: 11 Scheduled Re-acquisitions: 4 Successful: 4 2.2 FHST Updates: Scheduled: 25 Successful: 25 2.3 Operations Notes: Using ROP SR-3, the SSR pointers were set at 112/0959Z. The engineering status buffer limit for SESBSLD was adjusted at 112/1552Z and at 113/0345Z per ROP DF-18A. A 486 software dump was performed at 112/1754Z and at 112/1901Z, using ROP DF-01A. A TTR was generated when there was a required re-transmit during a NSSC-1 load at 112/2042Z. An operations request was utilized at 112/2113Z to transition the NCC PID control from the weighted neon temperatures to NICMOS dewar temperature. The primary PID set-point was changed from 73.0 degK to 74.8 degK and the NICMOS temperature valid flag in the NSSC-1 was set to indicate a valid dewar temperature, allowing the PID to actively manage the system using the primary control point. The ephemeris table was uplinked at 113/0143Z as directed by ROP DF-07A. The SSR-3 EDAC error counter was cleared at 113/0237Z per ROP DF-09A. HSTAR 8625 was written when the 486 status buffer issued a message indicating an FGS coarse mode angle check failure at 113/034137Z following a successful a primary acquisition at 13/033852Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of Servicing Mission Orbital Verification and the gradual resumption of normal science observations and calibrations. /CAW